# On compatibility of Koszul- and higher preprojective gradings

### v1

@@ Abstract.
We investigate compatibility of gradings for an almost Koszul or Koszul algebra $R$ that is also the higher preprojective algebra $\Pi_{n+1}(A)$ of an $n$-hereditary algebra $A$. For an $n$-representation finite algebra $A$, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with an almost Koszul grading. For a basic $n$-representation infinite algebra $A$ such that $\Pi_{n+1}(A)$ is graded coherent, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with a Koszul grading. From this we deduce that a higher preprojective grading of an (almost) Koszul algebra $R=\Pi_{n+1}(A)$ is, in both cases, isomorphic to a cut of the (almost) Koszul grading. Up to a further assumption on the tops of the degree 00 subalgebras for the different gradings, we also show a similar result without the basic assumption in the $n$-representation infinite case. As an application, we show that $n$-APR tilting preserves the property of being Koszul for $n$-representation infinite algebras that have graded coherent higher preprojective algebras.
@@ 1. Introduction
It is this interplay between the two gradings that we want to investigate in this article, both for the $n$-representation finite and the $n$-representation infinite case. The motivating question, which is made precise in 2.24, asks whether a given higher preprojective grading and a Koszul grading on some algebra $R$ need to be compatible in the way described above. We begin by explaining that this is not literally true and give some easy examples. However, the examples suggest that for two “incompatible” gradings we can find an automorphism that moves the higher preprojective grading to one that is compatible with the given Koszul grading. We give a proof of this statement in the $n$-representation finite case in 4.3. In the $n$-representation infinite case, we prove an analogous statement for basic algebras under a coherence assumption on the higher preprojective grading in 5.13, and a non-basic version over a fixed semisimple base ring in 5.14. The proof is based on the fact that if $A$ is ring-indecomposable and $n$-hereditary, and $\Pi_{n+1}(A)$ can be endowed with some (almost) Koszul grading, then $A$ can be endowed with a Koszul grading. In this way, our results can be seen as a converse to Grant-Iyama’s work, showing that $A$ can be given a Koszul grading if and only if $\Pi_{n+1}(A)$ can be given an (almost) Koszul grading. Additionally, this answers a question raised in [DG24a, Question 3.13] in the basic, $n$-representation tame case, which also applies to [DG24b]. In the $n$-representation tame case, our results have a clear geometric interpretation. The compatibility of gradings becomes equivalent to the commutation of two $k^{\ast}$-actions on the affine variety $\operatorname{Spec}(Z(\Pi_{n+1}(A)))$, as outlined in Remark 2.25. Our results also complement the work in [Thi20], where Thibault shows that certain Koszul algebras can not arise as higher preprojective algebras. Similarly, this shows that in order to construct preprojective gradings on Koszul quiver algebras, one can exhaust all gradings up to isomorphism by grading the quiver, as has been done in [HIO14, DG24a, DG24b, Gio19]. As an application, we show that $n$-APR tilting preserves Koszulity, and coherence of higher preprojective gradings.
@@ 2. Setup
We work over an algebraically closed field $k$ of characteristic zero.
@@ 2.3. $(n+1)$-preprojective algebras
In 5.13, we will additionally assume that the higher preprojective algebra ${}_{\bullet}R$ is graded coherent. This means that the category of finitely presented graded left ${}_{\bullet}R$-modules is closed under kernels. The coherence property and its implications for higher Auslander-Reiten theory have been studied in [Min12, MM11] as well as [HIO14, Section 6], and the question when the higher preprojective algebra is graded coherent was posed in [HIO14, Question 4.37]. We note that for example if $A$ is $n$-representation tame, then $\Pi_{n+1}(A)$ is graded coherent, and to our knowledge no example of a non-coherent higher preprojective grading is known.
@@ 2.5. Higher preprojective cuts
It was proven in [HIO14] that a ring-indecomposable $n$-hereditary algebra is either $n$-representation finite or $n$-representation infinite. This is reflected in the dimension of the higher preprojective algebra being finite or infinite. We therefore split the above question into the two cases of finite and infinite as well. We will give a positive answer to this question for the $n$-representation finite case. Under some mild coherence assumptions, we also prove this for the $n$-representation infinite case. The proof hinges on the fact that the preprojective degree 00 part can be given a Koszul grading, which may differ from the (almost) Koszul grading on the preprojective algebra.
@@ 3. Reduction to the graded basic case
Suppose $R_{\bullet}$ and ${}_{\bullet}R$ are a Koszul and a higher preprojective grading respectively, so that $A={}_{0}R$ is $n$-representation infinite. We may choose $e\in R_{0}$ for the reduction to the graded basic algebra $R_{\bullet}^{b}$. However, in general $e\not\in{}_{0}R$ need not be preprojectively homogeneous. This means that the subspaces $e({}_{i}R)e$ need not define a grading on $eRe$, so we can not transfer the higher preprojective grading this way. In order to remedy this, we set $f={}_{0}e$ and in this way obtain a full, homogeneous idempotent in ${}_{0}R$, which gives rise to a graded basic algebra $fRf=\bigoplus_{i\geq 0}(f({}_{i}R)f)$. However, since $R$ is infinite dimensional, the reductions $eRe$ and $fRf$ need not be isomorphic. We therefore will assume that $R_{0}\subseteq{}_{0}R$ in 5.14, so that the idempotent $e\in R_{0}$ is also homogeneous with respect to the preprojective grading. As pointed out in Example 2.23, this assumption is not fulfilled in general. In the spirit of the Wedderburn-Malcev theorem, one can hope that $R_{0}$ can be moved into ${}_{0}R$ by some automorphism, so we record the following question.
@@ 5.3. Nilpotency of generators
Let $A$ be $n$-hereditary, let $P$ be an indecomposable projective, and assume that $\operatorname{Hom}_{A}(D(A),A)=0$. Then any morphism from $P$ to $\tau_{n}^{-}(P)$ factors through some projective $A$-module $Q_{0}$ such that no summand of $Q_{0}$ is isomorphic to $P$.
@@
This space can be decomposed according to the decomposition of $A$ into indecomposable projectives, so we write
@@ 5.4. Koszulity of $n$-representation infinite algebras
In this section, we assume that $A={}_{0}R$ is $n$-representation infinite and that the graded simples of ${}_{\bullet}R$ and $R_{\bullet}$ are isomorphic as ungraded modules. As before, we assume that both gradings are locally finite dimensional, and additionally that ${}_{\bullet}R$ is graded coherent. Then we aim to show that $A$ is Koszul. We need the following notion of minimality for graded projective resolutions, which can be found in [MM11, Section 2].
@@
Let $R_{\bullet}$ be a locally finite Koszul grading and ${}_{\bullet}R$ be a graded coherent higher preprojective grading such that $A={}_{0}R$ is $n$-representation infinite. Suppose that the graded simple modules of $R_{\bullet}$ and ${}_{\bullet}R$ coincide as ungraded modules. Then $A$ is a Koszul algebra.
@@
is generated in cohomological degree 00 and $1$
Since ${}_{\bullet}R$ is assumed graded coherent and $S$ is finitely presented, we can find a graded free resolution $F^{\bullet}$ of $S$ so that each term $F^{i}$ is of finite rank. It then follows from the same argument as in [NVO04, Corollary 2.4.7] that we have
$\bigoplus_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle)\simeq\operatorname{Ext}^{i}_{R}(S,S).$
@@
5.5. Coherent preprojective gradings and cuts
@@
Let $R=R_{\bullet}$ be a graded basic Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists a coherent higher preprojective grading ${}_{\bullet}R$ on $R$, then there exists an automorphism that maps the higher preprojective grading to a higher preprojective cut. In particular, all coherent higher preprojective gradings are, up to isomorphism, higher preprojective cuts.
@@
Combining 5.10 and 5.6, we obtain that the graded simples for the Koszul grading and the higher preprojective grading coincide. Therefore, we can apply 5.12 to see that $A={}_{0}R$ is Koszul. Then, by 2.22, there exists a Koszul grading on $\Pi_{n+1}(A)\simeq R$ so that the higher preprojective grading on $\Pi_{n+1}(A)$ is a cut. Finally, by 2.9, there exists an automorphism taking the new Koszul grading on $\Pi_{n+1}(A)\simeq R$ to the given grading $R_{\bullet}$. Hence, this automorphism maps ${}_{\bullet}R$ to a cut. ∎
@@
Let $R=R_{\bullet}$ be a Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists a coherent higher preprojective grading ${}_{\bullet}R$ on $R$ such that $R_{0}\subseteq{}_{0}R$, then there exists a higher preprojective cut realising $R_{\bullet}=\Pi_{n+1}(A)$ so that $A\simeq{}_{0}R$.
@@
By assumption, we can reduce $R_{\bullet}$ to a graded basic algebra $R_{\bullet}^{b}$ by some homogeneous idempotent $e\in R_{0}\subseteq{}_{0}R$. Since $e$ is preprojectively homogeneous, it follows that $eRe$ inherits a higher preprojective grading. By 5.13, we obtain a higher preprojective cut, and by 3.8 this gives the higher preprojective cut on $R_{\bullet}$. ∎
We conclude with an application showing that $n$-APR tilting preserves being Koszul. For this, we need the following lemma showing that $n$-APR tilting preserves graded coherence of $(n+1)$-preprojective gradings. For the definition of $n$-APR tilting, we refer the reader to [IO11].
Lemma 5.15.
Let $A$ be $n$-representation infinite and $B$ be an $n$-APR tilt of $A$. Then $\Pi_{n+1}(A)$ is graded coherent if and only if $\Pi_{n+1}(B)$ is graded coherent.
@@
Since $B$ is an $n$-APR tilt of $A$, we have $\Pi_{n+1}(A)\simeq\Pi_{n+1}(B)$ as ungraded algebras. Furthermore, by the proof of [AO14, Theorem 5.6], $\Pi_{n+1}(A)$ and $\Pi_{n+1}(B)$ are graded equivalent in the sense of [GG82], and hence their categories of finitely presented modules are equivalent by [GG82, Theorem 5.4]; note that although [GG82] works with finitely generated modules, mutatis mutandis, the same arguments work for finitely presented modules. ∎
Corollary 5.16.
Let $A$ be basic and $n$-representation infinite, and assume that $\Pi_{n+1}(A)$ is graded coherent. Let $B$ be an $n$-APR tilt of $A$. Then $A$ is Koszul if and only if $B$ is Koszul.
Proof.
Note that since $B$ is an $n$-APR tilt of $A$, 5.15 implies that $\Pi_{n+1}(B)$ is graded coherent if $\Pi_{n+1}(A)$ is graded coherent. If $A$ is Koszul, then so is $\Pi_{n+1}(A)$ by 2.17, and the higher preprojective grading is a cut for this grading. Hence, the graded simple modules for the two gradings coincide as ungraded modules. Transferring the grading to $\Pi_{n+1}(B)$, we get a Koszul grading on $\Pi_{n+1}(B)$ satisfying the assumptions of 5.12, showing that $B$ is Koszul. ∎
@@ Acknowledgements
The proof of 5.15 was suggested to the authors by Steffen Oppermann. We thank Martin Herschend and Steffen Oppermann for many helpful discussions. The second author is grateful to have been supported by Norwegian Research Council project 301375, “Applications of reduction techniques and computations in representation theory”.

### v2

@@ Abstract.
We investigate compatibility of gradings for an almost Koszul or Koszul algebra $R$ that is also the higher preprojective algebra $\Pi_{n+1}(A)$ of an $n$-hereditary algebra $A$. For an $n$-representation finite algebra $A$, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with an almost Koszul grading. For a basic $n$-representation infinite algebra $A$, we show that $A$ must be Koszul if $\Pi_{n+1}(A)$ can be endowed with a Koszul grading. From this we deduce that a higher preprojective grading of an (almost) Koszul algebra $R=\Pi_{n+1}(A)$ is, in both cases, isomorphic to a cut of the (almost) Koszul grading. Up to a further assumption on the tops of the degree 00 subalgebras for the different gradings, we also show a similar result without the basic assumption in the $n$-representation infinite case. As an application, we show that $n$-APR tilting preserves the property of being Koszul for $n$-representation infinite algebras.
@@ 1. Introduction
It is this interplay between the two gradings that we want to investigate in this article, both for the $n$-representation finite and the $n$-representation infinite case. The motivating question, which is made precise in 2.24, asks whether a given higher preprojective grading and a Koszul grading on some algebra $R$ need to be compatible in the way described above. We begin by explaining that this is not literally true and give some easy examples. However, the examples suggest that for two “incompatible” gradings we can find an automorphism that moves the higher preprojective grading to one that is compatible with the given Koszul grading. We give a proof of this statement in the $n$-representation finite case in 4.3. In the $n$-representation infinite case, we prove an analogous statement for basic algebras in 5.14, and a non-basic version over a fixed semisimple base ring in 5.15. The proof is based on the fact that if $A$ is ring-indecomposable and $n$-hereditary, and $\Pi_{n+1}(A)$ can be endowed with some (almost) Koszul grading, then $A$ can be endowed with a Koszul grading. In this way, our results can be seen as a converse to Grant-Iyama’s work, showing that $A$ can be given a Koszul grading if and only if $\Pi_{n+1}(A)$ can be given an (almost) Koszul grading. Additionally, this answers a question raised in [DG24a, Question 3.13] in the basic, $n$-representation tame case, which also applies to [DG24b]. In the $n$-representation tame case, our results have a clear geometric interpretation. The compatibility of gradings becomes equivalent to the commutation of two $k^{\ast}$-actions on the affine variety $\operatorname{Spec}(Z(\Pi_{n+1}(A)))$, as outlined in Remark 2.25. Our results also complement the work in [Thi20], where Thibault shows that certain Koszul algebras can not arise as higher preprojective algebras. Similarly, this shows that in order to construct preprojective gradings on Koszul quiver algebras, one can exhaust all gradings up to isomorphism by grading the quiver, as has been done in [HIO14, DG24a, DG24b, Gio19]. As an application, we show that $n$-APR tilting preserves Koszulity.
@@ 2. Setup
We work over an algebraically closed field $k$ of characteristic zero, and we denote by $D({-})=\operatorname{Hom}_{k}({-},k)$ the usual $k$-duality.
@@ 2.3. $(n+1)$-preprojective algebras
An important property a higher preprojective algebra ${}_{\bullet}R$ may possess is that of being graded coherent. This means that the category of finitely presented graded left ${}_{\bullet}R$-modules is closed under kernels. The coherence property and its implications for higher Auslander-Reiten theory have been studied in [Min12, MM11] as well as [HIO14, Section 6], and the question when the higher preprojective algebra is graded coherent was posed in [HIO14, Question 4.37]. We note that for example if $A$ is $n$-representation tame, then $\Pi_{n+1}(A)$ is graded coherent. However, note that [Gel18] has used a relationship between certain kinds of $t$-structures on stable categories of maximal Cohen–Macaulay modules and the notion of absolute Koszulity to construct an infinite family of higher preprojective algebras that are not graded coherent. We therefore point out that we do not assume graded coherence of the higher preprojective algebra.
@@ 2.5. Higher preprojective cuts
It was proven in [HIO14] that a ring-indecomposable $n$-hereditary algebra is either $n$-representation finite or $n$-representation infinite. This is reflected in the dimension of the higher preprojective algebra being finite or infinite. We therefore split the above question into the two cases of finite and infinite as well. We will give a positive answer to this question for the $n$-representation finite case. We also prove this for basic algebras in the $n$-representation infinite case. The proof hinges on the fact that the preprojective degree 00 part can be given a Koszul grading, which may differ from the (almost) Koszul grading on the preprojective algebra.
@@ 3. Reduction to the graded basic case
Suppose $R_{\bullet}$ and ${}_{\bullet}R$ are a Koszul and a higher preprojective grading respectively, so that $A={}_{0}R$ is $n$-representation infinite. We may choose $e\in R_{0}$ for the reduction to the graded basic algebra $R_{\bullet}^{b}$. However, in general $e\not\in{}_{0}R$ need not be preprojectively homogeneous. This means that the subspaces $e({}_{i}R)e$ need not define a grading on $eRe$, so we can not transfer the higher preprojective grading this way. In order to remedy this, we set $f={}_{0}e$ and in this way obtain a full, homogeneous idempotent in ${}_{0}R$, which gives rise to a graded basic algebra $fRf=\bigoplus_{i\geq 0}(f({}_{i}R)f)$. However, since $R$ is infinite dimensional, the reductions $eRe$ and $fRf$ need not be isomorphic. We therefore will assume that $R_{0}\subseteq{}_{0}R$ in 5.15, so that the idempotent $e\in R_{0}$ is also homogeneous with respect to the preprojective grading. As pointed out in Example 2.23, this assumption is not fulfilled in general. In the spirit of the Wedderburn-Malcev theorem, one can hope that $R_{0}$ can be moved into ${}_{0}R$ by some automorphism, so we record the following question.
@@ 5.3. Nilpotency of generators
Let $A$ be $n$-hereditary, let $P$ be an indecomposable projective, and assume that $\operatorname{Hom}_{A}(D(A),A)=0$. Then any morphism from $P$ to $\tau_{n}^{-}(P)$ factors through some projective $A$-module $Q$ such that no summand of $Q$ is isomorphic to $P$.
@@
This space can be decomposed according to the decomposition of $A=\bigoplus_{i}P_{i}$ into indecomposable projectives, so we write
@@ 5.4. Koszulity of $n$-representation infinite algebras
In this section, we assume that $A={}_{0}R$ is $n$-representation infinite and that the graded simples of ${}_{\bullet}R$ and $R_{\bullet}$ are isomorphic as ungraded modules. As before, we assume that both gradings are locally finite dimensional. Then we aim to show that $A$ is Koszul. We need the following notion of minimality for graded projective resolutions, which can be found in [MM11, Section 2].
@@
The following lemma is a consequence of the fact that the gradings we consider are locally finite. The proof is almost verbatim that of [Mad11, Proposition 3.1.2].
Lemma 5.12.
Let $R_{\bullet}$ be a locally finite Koszul grading and ${}_{\bullet}R$ be a higher preprojective grading such that $A={}_{0}R$ is $n$-representation infinite, and write $S=\operatorname{top}(A)$. If $S\simeq R_{0}$ as ungraded $R$-modules, then
$\operatorname{Ext}^{i}_{R}(S,S)=\bigoplus_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle).$
Proof.
Since $S$ is concentrated in degree 00, it follows from the proof of [Mad11, Proposition 3.1.2] that
$\operatorname{Ext}_{R}^{i}(S,S)\simeq\prod_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle)$
for each $i\geq 0$. Since $R_{\bullet}$ is Koszul, we have that $\operatorname{Ext}^{0}_{R}(S,S)\simeq D(R_{0})$ and $\operatorname{Ext}^{1}_{R}(S,S)\simeq D(R_{1})$, which both are finite dimensional by assumption. Furthermore, since $R_{\bullet}$ is Koszul, we have that $\operatorname{Ext}^{\ast}_{R}(S,S)$ is generated in cohomological degrees 00 and $1$, so it follows that $\operatorname{Ext}_{R}^{i}(S,S)$ is finite dimensional for all $i\geq 0$. Therefore, each direct product $\prod_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle)$ is a finite direct product, and hence we obtain
$\operatorname{Ext}_{R}^{i}(S,S)\simeq\prod_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle)\simeq\bigoplus_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle).\qed$
@@
Let $R_{\bullet}$ be a locally finite Koszul grading and ${}_{\bullet}R$ be a higher preprojective grading such that $A={}_{0}R$ is $n$-representation infinite. Suppose that the graded simple modules of $R_{\bullet}$ and ${}_{\bullet}R$ coincide as ungraded modules. Then $A$ is a Koszul algebra.
@@
is generated in cohomological degree 00 and $1$. It follows from 5.12 that
$\operatorname{Ext}^{i}_{R}(S,S)\simeq\bigoplus_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle).$
@@
5.5. Preprojective gradings and cuts
@@
Let $R=R_{\bullet}$ be a graded basic Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists a higher preprojective grading ${}_{\bullet}R$ on $R$, then there exists an automorphism that maps the higher preprojective grading to a higher preprojective cut. In particular, all higher preprojective gradings are, up to isomorphism, higher preprojective cuts.
@@
Combining 5.10 and 5.6, we obtain that the graded simples for the Koszul grading and the higher preprojective grading coincide. Therefore, we can apply 5.13 to see that $A={}_{0}R$ is Koszul. Then, by 2.22, there exists a Koszul grading on $\Pi_{n+1}(A)\simeq R$ so that the higher preprojective grading on $\Pi_{n+1}(A)$ is a cut. Finally, by 2.9, there exists an automorphism taking the new Koszul grading on $\Pi_{n+1}(A)\simeq R$ to the given grading $R_{\bullet}$. Hence, this automorphism maps ${}_{\bullet}R$ to a cut. ∎
@@
Let $R=R_{\bullet}$ be a Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists a higher preprojective grading ${}_{\bullet}R$ on $R$ such that $R_{0}\subseteq{}_{0}R$, then there exists a higher preprojective cut realising $R_{\bullet}=\Pi_{n+1}(A)$ so that $A\simeq{}_{0}R$.
@@
By assumption, we can reduce $R_{\bullet}$ to a graded basic algebra $R_{\bullet}^{b}$ by some homogeneous idempotent $e\in R_{0}\subseteq{}_{0}R$. Since $e$ is preprojectively homogeneous, it follows that $eRe$ inherits a higher preprojective grading. By 5.14, we obtain a higher preprojective cut, and by 3.8 this gives the higher preprojective cut on $R_{\bullet}$. ∎
We conclude with an application showing that $n$-APR tilting preserves being Koszul. For the definition of $n$-APR tilting, we refer the reader to [IO11].
Corollary 5.16.
Let $A$ be basic and $n$-representation infinite and let $B$ be an $n$-APR tilt of $A$. Then $A$ is Koszul if and only if $B$ is Koszul.
@@
If $A$ is Koszul, then so is $\Pi_{n+1}(A)$ by 2.17, and the higher preprojective grading is a cut for this grading. Hence, the graded simple modules for the two gradings coincide as ungraded modules. Transferring the grading to $\Pi_{n+1}(B)$, we get a Koszul grading on $\Pi_{n+1}(B)$ satisfying the assumptions of 5.13, showing that $B$ is Koszul. ∎
@@ Acknowledgements
We thank Steffen Oppermann for helpful discussions, and for suggesting the proof of a Lemma which was used in a previous version of this article. We also thank Martin Herschend for many helpful discussions. The second author is grateful to have been supported by Norwegian Research Council project 301375, “Applications of reduction techniques and computations in representation theory”.